UV SPECTRUM AND PHOTOCHEMISTRY OF THE SIMPLEST CRIEGEE INTERMEDIATE CH_{2}OO

JOSEPH M. BEAMES, \underline{FANG LIU}, LU LU, and MARSHA I. LESTER, Department of Chemsitry, University of Pennsylvania, Philadelphia, PA 19104-6323.

In the troposphere, ozonolysis of ethene as well as other terminal alkenes, such as isoprene, results in formation of the simplest Criegee intermediate, CH_{2}OO, which gives rise to many products of atmospheric significance. In the laboratory, CH_{2}OO is prepared in a quartz capillary tube reactor using a new synthetic route^{a} based on 248 nm photolysis of CH_{2}I_{2} and subsequent reaction with O_{2}. The CH_{2}OO and other products undergo supersonic expansion, and are detected using fixed frequency VUV ionization (118 nm) with mass (m/z=46) and isomer selectivity. Spectroscopic studies of CH_{2}OO are carried out by UV laser excitation on the B ^{1}A′-X ^{1}A′ transition prior to photoionization, which results in significant depletion of the ion signal (approaching 100%) near the peak of a broad absorption profile centered at 335 nm.^{b} The large depletion and broad absorption are indicative of rapid dissociation, consistent with the repulsive B ^{1}A′ state potential in the O-O coordinate that has been computed theoretically. The experimental spectrum is in very good agreement with the absorption spectrum calculated using the one-dimensional reflection principle. The atmospheric lifetime of CH_{2}OO due to solar photolysis at midday is estimated to be ∼ 1 s. These findings provide a UV spectral signature of the strong π^{∗}-π transition associated with the four π electrons on the carbonyl oxide group in CH_{2}OO. Extensions of this work to methyl substituted Criegee intermediates reveal similar absorption features arising from the COO π -system .

aO. Welz, J. D. Savee, D. L. Osborn, S. S. Vasu, C. J. Percival, D. E. Shallcross, and C. A. Taatjes, Science 335, 204 (2012).

bJ. M. Beames, F. Liu, L. Lu, and M. I. Lester, J. Am. Chem. Soc. 134, 20045 (2012).